# ExtractBench: A Benchmark and Evaluation Methodology for Complex Structured Extraction

### v1

@@ 2.1. Design Philosophy
10-K/QCreditResearchResumesSports01010202030304040Avg. Tokens (k)Input vs Output TokensInput (PDF)Output (JSON)0202040406060808010-K/QCreditResearchResumesSportsCompression RatioInput/Output Ratio
@@
Figure 2 reveals two contrasting task types: credit agreements are “needle-in-haystack” tasks requiring models to locate sparse fields in lengthy legal prose, while research papers are “enumeration” tasks where citation arrays expand output beyond input volume. This explains why credit agreements – despite being the longest documents – achieve the highest pass rate: output volume, not input length, is the binding constraint.
@@ Appendix B Additional Tables and Figures
Tables 913 provide per-domain breakdowns of model performance, complementing the aggregate results in Table 5.

### v2

@@ 2.1. Design Philosophy
10-K/QCreditResearchResumesSports1k10k100kAvg. Tokens (k)Input vs Output TokensInput (Vision)Output (JSON)020020040040060060010-K/QCreditResearchResumesSportsCompression RatioInput/Output Ratio
@@
[…]
@@ Appendix B Additional Tables and Figures
Tables 913 provide per-domain breakdowns of model performance, complementing the aggregate results in Table 5. Table 14 provides per-document token statistics.
@@
Table 14. Per-document token statistics. Input tokens are Vision tokens (raw image representation). Compression ratio = Input $\div$ Output.
Domain Document Pages Input (Vision) Output Ratio
Sports Results ma_2023_sw_M-table1 2 4,634 468 9.9
ma_2023_sw_M-table2 3 6,951 250 27.8
ma_2023_sw_M-table3 3 6,951 268 25.9
ma_2023_sw_M-table4 3 6,951 385 18.1
ma_2023_sw_M-table5 4 9,268 551 16.8
Research Papers NIPS-1989 (backprop) 9 22,810 815 28.0
FlashAttention-3 22 55,726 4,245 13.1
Dimensionality reduction survey 35 81,095 7,551 10.7
RAG survey 18 45,594 13,403 3.4
VLM survey 22 55,726 20,766 2.7
LLM survey 144 364,752 91,261 4.0
SEC 10-K/10Q ADP 10-Q FY25Q2 43 99,588 2,734 36.4
Nike 10-Q FY25Q2 48 111,168 3,045 36.5
THO 10-Q FY25Q2 53 122,748 2,779 44.2
McKesson 10-Q FY25Q2 58 134,328 2,824 47.6
WDC 10-Q FY25Q2 59 136,644 3,003 45.5
Cisco 10-Q FY25Q2 79 182,964 3,137 58.3
Dell 10-Q FY25Q2 82 189,912 2,538 74.8
Credit Agreements Adobe (2000) 97 245,701 306 803
Disney (2022) 102 236,232 510 463
Boeing (2003) 110 278,630 360 774
3M (2019) 112 283,696 654 434
Expel (2023) 123 284,868 365 780
Cisco (2007) 126 319,158 725 440
BKRF (2020) 140 324,240 551 588
Trimble (2022) 143 331,188 434 763
IBM (2019) 197 499,001 798 625
Amazon (2014) 218 552,194 480 1,150
Resumes Resume-Finance 1 2,533 432 5.9
Resume-Legal 1 2,533 521 4.9
Resume-Med 1 2,533 680 3.7
Resume-IT 1 2,533 388 6.5
Resume-Marketing 3 7,599 861 8.8
Resume-Academic01 7 17,731 2,996 5.9
Resume-Academic02 7 17,731 2,861 6.2